### v1

@@ 1 Introduction
Neutrinos reaching the SND@LHC detector are energetic enough to produce heavy quarks, such as the charm quark, in the final state. This allows SND@LHC to be a laboratory for testing LUV effects in quark-neutrino scattering processes involving a heavy quark. The SND@LHC detector also aims at exploring the possibility of detecting new particles that scatter similar to neutrinos, such as light dark matter (LDM) particles which interact with SM particles through portal mediators. This study focuses on estimating the sensitivity of the SND@LHC detector in detecting new-physics phenomena, particularly in terms of effective four-fermion operators.
@@
A key feature of SND@LHC is its high efficiency in identifying neutrino flavors. Electron neutrinos can be identified via their electromagnetic showers with 99% efficiency. Charged-current muon neutrino interactions can be detected with 69% efficiency by requiring the presence of at least one muon track. Neutral current interactions are correctly tagged with 99% efficiency. Thanks to the micrometric resolution of the ECC, tau neutrinos can be identified via a decay vertex that is displaced from the primary interaction point. The $\nu_{\tau}$ detection efficiency for SND@LHC ranges from 48% to 54% depending on the decay mode of the $\tau$ lepton Ahdida:2750060 .
In this work, we will explore LUV effects in the scattering process $\nu_{\ell}+q\to c+\ell^{-}$ at SND@LHC, where $q$ represents down-type quarks in the target (tungsten) nuclei. The incoming neutrinos are highly energetic so deep-inelastic scattering (DIS) is the underlying process. In the SM, the contributions from a $b$-quark in the initial state are suppressed due to the relative smallness of the Cabibbo-Kobayashi-Maskawa (CKM) matrix element $V_{cb}$ compared to contributions from a $d$ or $s$ quark in the initial state. For the same reason, the scattering process where an up-type quark from the target nuclei contributes in the initial state to produce a $b$ quark in the final state is also suppressed and is not considered here. We adopt an effective field theory (EFT) framework starting with all dimension 6 operators that can contribute to the process. We then constrain the Wilson Coefficients (WCs) of the effective operators using low-energy measurements of two- and three-body $D$ decays. While constraints on NP from semileptonic charmed mesons have been considered before Becirevic:2020rzi , here we present a more exhaustive fit to the low-energy charm semileptonic decays. In our low-energy observables, we consider both two-body leptonic and three-body semileptonic decays of charmed mesons where we include decays not only to pseudoscalar final states but also to vector final states and include in our fits decays to the $\eta$ and $\eta^{\prime}$ mesons. We also consider constraints from charmed baryon decays. After the low energy fit, various benchmark values for the WCs are then used to study the prospects of detecting NP at SND@LHC.
@@ 3.3 Three-body baryonic decays
Table 1: List of observables used to constrain the new physics coefficients along with their SM predictions and measured branching fractions. The measured values are taken from PDG ParticleDataGroup:2022pth while the SM predictions are computed using the form factors defined in Appendix C and CKM elements listed in Table 8.
@@ 3.4 Low energy fit results
We fit the new physics Wilson coefficients in Eq. (4) to the observables listed in Table 1 separately for the electron and muon modes. In Table 1, we also list the SM expectation and the measured values of the branching fractions for each decay mode. For the numerical analysis we consider the most updated value for the decay constants of the D mesons for the leptonic decays from Ref. FlavourLatticeAveragingGroupFLAG:2021npn . The form factors for the semileptonic charm decays are not much developed. In any case, we use the available results on the form factors from references listed in Appendix C. Note that there are 12 measurements for the electron modes and 14 for the muon modes which are utilized to fit to the respective flavour specific Wilson coefficients. For the tau mode, however, we only have measurements for the two body leptonic decay of the $D_{(s)}$ meson. Furthermore, since the $\nu_{\tau}$ detection efficiency at SND is low compared to the other lepton flavour, it is challenging to obtain any meaningful constraint on the new physics parameters. Hence, we do not fit to tau measurements in our analysis.
@@ 4 Deep Inelastic Neutrino-Nucleon Scattering
$\chi^{2}\,=\,\sum_{i=1}^{n}\left(\frac{N_{i}^{\rm NP+SM}-N_{i}^{\rm SM}}{\sqrt{N_{i}^{\rm SM+NP}}}\right)^{2}$ (22)
@@
$\chi^{2}=\sum_{i=1}^{n}\left(\frac{R_{\mu,e}^{\rm NP+SM}-R_{\mu,e}^{\rm SM}}{\sigma_{R_{\mu,e}^{\rm SM+NP}}}\right)^{2},$ (23)
@@ 5 Results
In principle, NP affecting charm quarks, as encoded in the effective Hamiltonian of Eq. (4), will also impact the production flux of neutrinos. Ref. Kling:2021gos demonstrates that charm-hadron decays predominantly drive the electron-neutrino flux at energies above 200 GeV, while they contribute equally to the muon-neutrino flux, along with pion decays, at energies above 400 GeV. These effects have been thoroughly analyzed in Ref. Falkowski:2021bkq for all neutrino flavors. Their study shows that an NP pseudoscalar interaction with order-one coupling significantly amplifies the flux at production due to $D_{s}$ decays, increasing the electron-neutrino flux by up to $10^{3}\times$ and the muon-neutrino flux by up to 27×27$\times$27 × that in the SM. Similar to Ref. Falkowski:2021bkq , we assume that among all charm-hadron decays, only 2-body $D_{s}$ decays contribute significantly. Given the significant enhancement observed for the pseudoscalar coupling, we compute the energy distributions by accounting for the modifications in neutrino flux during production, as detailed in Ref. Falkowski:2021bkq . The resulting distributions are presented in Fig. 3 for the electron neutrino and Fig. 4 for the muon neutrino.
@@
Figure 3: Number of signal (SM + NP, dashed) and background (SM only, solid) $\nu_{e}+N\to e+X$ events as functions of the incoming neutrino energy, $E_{\nu}$. The integrated luminosity ${\cal L}$ has been set to the SND@LHC current projection of 290 fb-1 for the lower (purple) histograms, and to its high-lumi upgrade projection of 3000 fb-1 for the upper (pink) histograms. These results use $|{\rm Re}[\epsilon_{P}]|=0.002$ of model 13 in Table 2 and the SND@LHC detector pseudorapidity range of $8.47.2\,<\eta\,<$ 8.4. The left plot assumes no charm tagging, while the right plot assumes 100% charm tagging.
[…]
Figure 4: Number of signal (SM + NP, dashed) and background (SM only, solid) $\nu_{\mu}+N\to\mu+X$ events as functions of the incoming neutrino energy, $E_{\nu}$. The integrated luminosity ${\cal L}$ has been set to the SND@LHC current projection of 290 fb-1 for the lower (dark blue) histograms, and to its high-lumi upgrade projection of 3000 fb-1 for the upper (light blue) histograms. These results use $\epsilon_{P}=-0.007+0.011i$ of model 17 in Table 3 and the SND@LHC detector pseudorapidity range of $8.47.2\,<\eta\,<$ 8.4. The left plot assumes no charm tagging, while the right plot assumes 100% charm tagging.
As shown in Fig. 3, the background dominates at energies below 200 GeV for the electron case. Therefore, we perform a $\chi^{2}$ analysis for bins with $E_{\nu}>200$ GeV. The sensitivity results focusing on electrons in the final state are presented in Table 4. Only those cases where the sensitivity exceeds $0.5\,\sigma$ have been included. Due to the flux enhancement from pseudoscalar NP interactions, only the cases with significant pseudoscalar couplings have considerable significance. While the electron-neutrino flux is nearly an order of magnitude smaller than the muon-neutrino flux for $E_{\nu}>100$ GeV (see Fig. 2), this difference is compensated for by the enhancement in production associated with pseudoscalar NP interactions. Consequently, the sensitivity remains high in scenarios with electrons in the final state. As discussed in Section 3.4 and shown in Table 2, the $\epsilon_{P}$ coupling for cases involving electrons, remain largely unconstrained in many models preventing sensitivity calculations for these specific cases. Nonetheless, a luminosity of 290 fb-1 at SND@LHC allows us to constrain an $\epsilon_{P}$ value of $10^{-3}$ at a $2\,\sigma$ significance level.
Models Sens($\nu_{e}+N\to e+X$) Sens$\left(\displaystyle\frac{\nu_{\mu}+N\to\mu+X}{\nu_{e^{-}}+N\to e^{-}+X}\right)$
[…]
The sensitivity results for muons in the final state are presented in Table 5. Here, the sensitivity is influenced not only by the pseudoscalar coupling, but also, albeit to a lesser extent, by the right-handed vector coupling. The sensitivity results listed in Table 5 were evaluated based on the central values of NP couplings listed in Table 3.
Models Sens($\nu_{\mu}+N\to\mu+X$) Sens$\left(\displaystyle\frac{\nu_{\mu}+N\to\mu+X}{\nu_{e^{-}}+N\to e^{-}+X}\right)$
@@
Notably, the high-luminosity upgrade of SND@LHC can collect ten times more data Abbaneo:2895224 , significantly enhancing sensitivity to NP, even in models where no significant sensitivity is observed at = $\mathcal{L}\,=\,290\,{\rm fb}^{-1}$. We assess the sensitivity across different lepton flavors for a luminosity of 3000 fb-1, as shown in Table 6. Figs. 3 and 4 also include event distributions for the high-lumi upgrade of SND@LHC.
Models Sens($\nu_{\mu}+N\to\mu+X$) Sens($\nu_{e}+N\to e+X$)
@@
Table 6: Sensitivity of detecting NP in the signal $\nu_{\mu}+N\to\mu+X$ and $\nu_{e}+N\to e+X$ at SND@LHC, based on models listed in Table 2 and Table 3, assuming a luminosity of 3000 fb-1.
6 Conclusion
[…]
Acknowledgments: This work was financially supported by the U.S. National Science Foundation under Grant No. PHY-2310627 (BB) and PHY-2309937 (AD). BB acknowledges support in part from the NSF Grant No. PHY-2309135 to the Kavli Institute of Theoretical Physics (KITP) where part of this work was completed. DS is supported by funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 101002846 (ERC CoG “CosmoChart”). The work of EG is supported by the Swiss National Science Foundation (SNSF) under grant number 202065 (“Ambizione”). We are grateful to the SND@LHC collaboration for several useful discussions.

### v2

@@ 1 Introduction
Neutrinos reaching the SND@LHC detector are energetic enough to produce heavy quarks, such as the charm quark, in the final state. This allows SND@LHC to be a laboratory for testing LUV effects in quark-neutrino scattering processes involving a heavy quark. The SND@LHC detector also aims at exploring the possibility of detecting new particles that scatter similar to neutrinos, such as light dark matter (LDM) particles which interact with SM particles through portal mediators. This study focuses on estimating the sensitivity of the SND@LHC detector in detecting new-physics phenomena, considering various benchmark new-physics couplings derived from low-energy fits.
@@
A key feature of SND@LHC is its high efficiency in identifying neutrino flavors. Electron neutrinos can be identified via their electromagnetic showers with 99% efficiency. Charged-current muon neutrino interactions can be detected with 69% efficiency by requiring the presence of at least one muon track. Neutral current interactions are correctly tagged with 99% efficiency. Thanks to the micrometric resolution of the ECC, tau neutrinos can be identified via a decay vertex that is displaced from the primary interaction point. The $\nu_{\tau}$ detection efficiency for SND@LHC ranges from 48% to 54% depending on the decay mode of the $\tau$ lepton Ahdida:2750060 ; SNDLHC:2022ihg .
[…]
@@ 3.3 Three-body baryonic decays
Table 1: List of observables used to constrain the new physics coefficients along with their SM predictions and measured branching fractions. The measured values are taken from the PDG ParticleDataGroup:2022pth while the SM predictions are computed using the form factors defined in Appendix C and the CKM matrix elements listed in Table 7.
@@ 3.4 Low energy fit results
We fit the new physics Wilson coefficients in Eq. (4) to the observables listed in Table 1 separately for the electron and muon modes. In Table 1, we list the SM expectation and the measured values of the branching fractions for each decay mode. For our numerical analysis, we take the values of $D$-meson decay constants and lattice determinations of relevant CKM matrix elements from the FLAG Review FlavourLatticeAveragingGroupFLAG:2021npn  [^1] Ref. Bolognani:2024cmr points out that the inclusion of a universal electroweak correction in the Wilson coefficient $\epsilon_{L}$ leads to a $\sim$ 1% increase in $\epsilon_{L}$. This also leads to a smaller value of $|V_{cs}|$ compared to the one obtained by the PDG ParticleDataGroup:2022pth . In this work, however, we use $\epsilon_{L}=1$ and $|V_{cs}|$ as given in the FLAG Review FlavourLatticeAveragingGroupFLAG:2021npn .. In Appendix C we list the relevant numerical inputs and form factors used in this article. Note that there are 12 measurements for the electron modes and 14 for the muon modes which are utilized to fit to the respective flavour specific Wilson coefficients. For the tau mode, however, we only have measurements for the two body leptonic decay of the $D_{(s)}$ meson. Furthermore, since the $\nu_{\tau}$ detection efficiency at SND is low compared to the other lepton flavour, it is challenging to obtain any meaningful constraint on the new physics parameters. Hence, we do not fit to tau measurements in our analysis.
@@ 4 Deep Inelastic Neutrino-Nucleon Scattering
$\chi^{2}\,=\,\sum_{i=1}^{n}\left(\frac{N_{i}^{\rm SM+NP}-N_{i}^{\rm SM}}{\sqrt{N_{i}^{\rm SM+NP}}}\right)^{2}$ (22)
@@
$\chi^{2}=\sum_{i=1}^{n}\left(\frac{R_{\mu,e}^{\rm SM+NP}-R_{\mu,e}^{\rm SM}}{\sigma_{R_{\mu,e}^{\rm SM+NP}}}\right)^{2},$ (23)
@@ 5 Results
In principle, NP affecting charm quarks, as encoded in the effective Hamiltonian of Eq. (4), will also impact the production flux of neutrinos. Ref. Kling:2021gos demonstrates that charm-hadron decays predominantly drive the electron-neutrino flux at energies above 200 GeV, while they contribute equally to the muon-neutrino flux, along with pion decays, at energies above 400 GeV. These effects have been thoroughly analyzed in Ref. Falkowski:2021bkq for all neutrino flavors. Their study shows that an NP pseudoscalar interaction with order-one coupling significantly amplifies the flux at production due to 2-body $D_{s}$ decays, increasing the electron-neutrino flux by up to $10^{3}\times$ and the muon-neutrino flux by up to 27×27$\times$27 × than in the SM. However, $\nu_{e}$ and $\nu_{\mu}$ production are dominated by three-body $D_{s}$ decays, which unlike two-body $D_{s}$ decays are not similarly enhanced. Therefore, these production and detection enhancements do not impact our analysis, as we only consider NP in the charm sector. We emphasize that, even if NP were present in first-generation quarks, stringent low-energy constraints impose strong limits on the corresponding Wilson coefficients, rendering current neutrino detectors ineffective in probing such NP operators. Furthermore, in our analysis even at a high luminosity, significant sensitivity to NP affecting muons is only observed when the final state is charm tagged. We, therefore, assume a 100% charm-tagging efficiency.
@@
Figure 3: Number of signal (SM + NP, dashed) and background (SM only, solid) events as functions of the incoming neutrino energy ($E_{\nu}$) for $\nu_{e}+N\to e^{-}+X_{c}$ (left plot) and $\nu_{\mu}+N\to\mu+X_{c}$ (right plot.) The integrated luminosity ${\cal L}$ has been set to the SND@LHC current projection of 290 fb-1 for the lower (dark blue/dark pink) histograms, and to its high-lumi upgrade projection of 3000 fb-1 for the upper (light blue/light pink) histograms. These results use couplings of model 17 from Tables 2 and 3, alongwith the pseudorapidity range $8.47.2\,<\eta\,<$ 8.4 corresponding to the SND@LHC detector.
The resulting event distributions for the electron and muon neutrinos are presented in Fig. 3. As can be seen in Fig. 3, the NP contribution is negligible for the electron neutrino. In Table 4, we present our results for the sensitivities of detecting NP with muons in the final state. This sensitivity is influenced by the left- and right-handed vector couplings. These results are based on the central values of NP couplings listed in Table 3. Here, we only include those cases where the sensitivity exceeds $0.1\,\sigma$.
Models Sens($\nu_{\mu}+N\to\mu+X_{c}$) Sens$\left(\displaystyle\frac{\nu_{\mu}+N\to\mu+X_{c}}{\nu_{e^{-}}+N\to e^{-}+X_{c}}\right)$
@@
The high-luminosity upgrade of SND@LHC can collect ten times more data Abbaneo:2895224 , significantly enhancing the sensitivity to NP, even in models where no significant sensitivity is observed at = $\mathcal{L}\,=\,290\,{\rm fb}^{-1}$. We assess the sensitivity for a luminosity of 3000 fb-1 in Table 5. Fig. 3 also includes event distributions for the high-lumi upgrade of SND@LHC.
Models Sens($\nu_{\mu}+N\to\mu+X_{c}$) Sens$\left(\displaystyle\frac{\nu_{\mu}+N\to\mu+X_{c}}{\nu_{e^{-}}+N\to e^{-}+X_{c}}\right)$
@@
Table 5: Sensitivity of detecting NP in the signal $\nu_{\mu}+N\to\mu+X_{c}$ at SND@LHC, based on models listed in Table 3, assuming a luminosity of 3000 fb-1.
Finally, we note the importance of $\nu_{\tau}$ scattering at experiments such as SND@LHC. NP effective operators involving both the charm quark and the $\tau$ neutrino are difficult to constrain from low-energy measurements. Due to the relative closeness in the $\tau$ and charm masses, three-body decays are suppressed or prohibited due to phase-space effects. In these cases, neutrino scattering experiments can shed a significant amount of light on such NP operators. In this article, however, we restrict ourselves to the cases where neutrino scattering experiments can provide information complementary to low-energy constraints, and as such do not consider $\tau$-neutrino scattering.
6 Conclusion
In this paper, we have presented a detailed sensitivity analysis of new physics (NP) and lepton universality violation (LUV) in neutrino scattering at SND@LHC, focusing on processes involving a charm quark in the final state. Within the framework of effective field theory, NP effects were described using higher-dimensional four-Fermi operators. The Wilson coefficients for these dimension-6 operators were determined from fits to low-energy data, particularly from decays of charmed mesons and baryons. Using benchmark NP points derived from these low-energy analyses, we provide sensitivity estimates for detecting NP and LUV effects at SND@LHC. Both production and scattering processes of neutrinos were considered in evaluating NP effects.
@@ 6 Conclusion
We also believe that a more comprehensive, multi-dimensional analysis could further improve the sensitivity. Additionally, our findings indicate that these detectors may be sensitive to NP operators that evade low-energy constraints but become relevant at high-energy scales. One such operator, as proposed in Ref. Datta:1996gg , fits this profile, though we leave a detailed investigation for future work. Furthermore, since effective couplings involving charmed particle decays with a final state $\nu_{\tau}$ receive less stringent constraints from low-energy experiments due to phase space restrictions, we note that neutrino scattering experiments can provide useful information about NP in such effective operators.
Acknowledgments: This work was financially supported by the U.S. National Science Foundation under Grant No. PHY-2310627 (BB) and PHY-2309937 (AD). BB acknowledges support in part from the NSF Grant No. PHY-2309135 to the Kavli Institute of Theoretical Physics (KITP) where part of this work was completed. DS is supported by funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 101002846 (ERC CoG “CosmoChart”). The work of EG is supported by the Swiss National Science Foundation (SNSF) under grant number 202065 (“Ambizione”). We are grateful to the SND@LHC collaboration for several useful discussions. We thank J. Kopp, F. Kling, and Z. Tabrizi for pointing out that three-body decays play a vital role in the production of electron and muon neutrinos for SND@LHC.